OpenStudy (anonymous):

Simplify the radical: square root 200

OpenStudy (owlcoffee):

\[\sqrt{200}\] By simple multiplication we can do this: \[\sqrt{(2)(100)}\] And applying some radication properties, we get: \[(\sqrt{2})(\sqrt{100})\] And we know that Sqrt of 100 is 10: \[(\sqrt{2})(10)\] And there we go, it's simplified :)

OpenStudy (anonymous):

oh, thank you :)